Two stars differ by 8 absolute magnitude.what is the ratio of there luminosities?


1
Expert's answer
2021-01-05T12:04:57-0500

For Pogson's formula, we write

"\frac{E_1}{E_2}=2.512^{M_2-M_1}=2.512^8=1585"